How fast can a raindrop (r=1.5mm) fall through air if the flow around it is to be done close to the turbulent, i.e., for Reynolds Number close to 10? for air, viscosity= 1.8 x10^-5 Pa.s and flowing with speed of 1.29kg/m³ answer?
1
Expert's answer
2020-11-16T06:36:29-0500
"Re=\frac{\rho u d}{\mu}\\10=\frac{(1.29)(0.003)u}{1.85\cdot 10^{-5}}\\u=0.048\frac{m}{s}"